OpenStudy (anonymous):

|2(x - 3) + 5| - 4 = 6x

OpenStudy (hoblos):

|2(x - 3) + 5| - 4 = 6x |2x-6+5| = 6x+4 |2x-1| = 6x+4 2x-1 = 6x+4 or 2x-1=-6x-4 4x=-5 8x=-3 x=-5/4 x=-3/8

OpenStudy (mertsj):

-5/4 is not correct

OpenStudy (hoblos):

why?!

OpenStudy (anonymous):

it was -3 and 8

OpenStudy (mertsj):

It is incorrect because it does not make the original equation true.